# CryoFastAR: Fast Cryo-EM Ab initio Reconstruction Made Easy

### v1

@@ 6 Discussion
As the first geometric foundation model for direct pose estimation from multi-view noisy cryo-EM images, our model dose have several limitations. First, it is trained primarily on a large-scale simulated dataset with limited real-world data, resulting in performance degradation on unseen real data due to the domain gap and the scarcity of annotated experimental data. This issue could be mitigated through more realistic cryo-EM image synthesis [51] or by collecting accurately labeled real-world data. Second, while our model achieves state-of-the-art performance on simulated datasets, it processes only a limited subset of available images in a single forward pass, constraining final reconstruction accuracy. Recent methods like Fast3R [50] and Spann3R [44], which employ advanced decoders and inference strategies to handle larger image batches, providing promising directions for improving scalability. Lastly, our model does not support heterogeneous reconstruction, whereas recent work such as CUT3R [45] explores promising approaches for dynamic scene reconstruction.
@@ Conclusion.
We have introduced CryoFastAR, the first geometric foundation model for fast ab initio reconstruction in cryo-EM. By leveraging a vision transformer-based architecture and predicting Fourier planar maps, CryoFastAR has effectively integrated multi-view information without requiring iterative pose search or precomputed CTF parameters. Extensive experiments have demonstrated that our method achieves competitive performance on real and synthetic datasets while significantly reducing computational costs. Our findings have highlighted the potential of feed-forward architectures in cryo-EM reconstruction, paving the way for more efficient and scalable structural analysis. We will release our code, models, and datasets to stimulate further research.
@@ B Details of Dataset
We generate 3 simulated datasets for evaluation of the baselines. Each dataset has the same simulation procedure: 1) generating 50000 particles with uniformly sampled rotations, 2) adding CTF corruption 3) translating image in $[-10,10]$ pixels along $x$ and $y$ axes, respectively, 4) add Gaussian noise to adjust the signal-to-noise ratio (SNR) of the image to 0.1. Each structure in the dataset has a different spatial resolution in terms of Å per pixel(Apix) when the PDB [3] structure is converted into a volume density map using EMAN2 [41]. The PDB ID and Apix for each dataset is as follows:
@@ B.3 Contrast Transfer Function
$\text{CTF}(w,\lambda,\textbf{g},\Delta f,C_{s},\Delta\varphi)=-\sqrt{1-w^{2}}\sin[\chi(\lambda,|\textbf{g}|,\Delta f,C_{s},\Delta\varphi)]-w\cos[\chi(\lambda,|\textbf{g}|,\Delta f,C_{s},\Delta\varphi)],$ (14)
@@ E Details of Simulated Training Dataset Construction
We begin with the 113,600 curated 3D atomic structures and generate noisy projection images for training through the following steps: For each 3D structure, we convert it into a $128^{3}$ 3D volume density map using EMAN2 [41], the pixel size is set by the maximum length of the structure to make sure the whole structure is inside. Then we uniformly sample 100 projection images in SO(3) space, representing different views of the volume. Each clean projection image is then modified by randomly applying a Contrast Transfer Function (CTF) sampled from real distribution including 152,385 CTF parameters as described in  [39], simulating imaging system imperfections such as aberrations or blurring. A random 2D shift, within the range of $[-10,10]^{2}$, is applied to each image to account for potential translational errors during imaging. Finally, Gaussian noise is added to each image to ensure the signal-to-noise ratio (SNR) matches the desired level, simulating the noise characteristics in experimental data. This augmentation pipeline produces diverse, realistic projections for model training.

### v2

@@ 6 Discussion
As the first geometric foundation model for direct pose estimation from multi-view noisy cryo-EM images, our method has some limitations. It is mainly trained on simulated data, leading to performance drops on real images due to domain gaps and limited annotations. This can be mitigated with more realistic simulations [51] or high-quality labeled data. Additionally, it processes only a subset of images per forward pass, limiting reconstruction accuracy. Scalable approaches like Fast3R [50] and Spann3R [45] offer promising solutions.
@@ Conclusion.
We have introduced CryoFastAR, the first geometric foundation model for fast ab initio reconstruction in cryo-EM. By leveraging a ViT-based architecture and predicting Fourier planar maps, CryoFastAR has effectively integrated multi-view information without requiring iterative pose search. Extensive experiments have demonstrated that our method achieves competitive performance on real and synthetic datasets while significantly reducing computational costs. Our findings have highlighted the potential of feed-forward architectures in cryo-EM reconstruction, paving the way for more efficient and scalable structural analysis.
7 Acknowledgment
This work was supported by ShanghaiTech University’s HPC Platform. We would like to thank the Cellverse team for their valuable discussions.
—Supplementary Material—
CryoFastAR: Fast Cryo-EM Ab initio Reconstruction Made Easy
Figure 6: Qualitative Result. We compare our visual quality with all other baselines before and after the refinement for Spliceosome’s simulated dataset. The results show that our method is comparable to them before refinement and achieves the best performance after the refinement.
@@ B Details of Dataset
We generate 3 simulated datasets for evaluation of the baselines. Each dataset has the same simulation procedure: 1) generating 50000 particles with uniformly sampled rotations, 2) adding CTF corruption 3) translating image in [10,10][-10,10] pixels along $x$ and $y$ axes, respectively, 4) add Gaussian noise to adjust the signal-to-noise ratio (SNR) of the image to 0.1. Each structure in the dataset has a different spatial resolution in terms of Å per pixel(Apix) when the PDB [3] structure is converted into a volume density map using EMAN2 [42]. The PDB ID and Apix for each dataset is as follows:
@@ B.3 Contrast Transfer Function
$\text{CTF}(w,\lambda,\textbf{g},\Delta f,C_{s},\Delta\varphi)={}$ (14)
$-\sqrt{1-w^{2}}\sin[\chi(\lambda,|\textbf{g}|,\Delta f,C_{s},\Delta\varphi)]$
$-w\cos[\chi(\lambda,|\textbf{g}|,\Delta f,C_{s},\Delta\varphi)]$
@@ E Details of Simulated Training Dataset Construction
We begin with the 113,600 curated 3D atomic structures and generate noisy projection images for training through the following steps: For each 3D structure, we convert it into a $128^{3}$ 3D volume density map using EMAN2 [42], the pixel size is set by the maximum length of the structure to make sure the whole structure is inside. Then we uniformly sample 100 projection images in SO(3) space, representing different views of the volume. Each clean projection image is then modified by randomly applying a Contrast Transfer Function (CTF) sampled from real distribution including 152,385 CTF parameters as described in  [40], simulating imaging system imperfections such as aberrations or blurring. A random 2D shift, within the range of $10,10]^{2}$, is applied to each image to account for potential translational errors during imaging. Finally, Gaussian noise is added to each image to ensure the signal-to-noise ratio (SNR) matches the desired level, simulating the noise characteristics in experimental data. This augmentation pipeline produces diverse, realistic projections for model training.